# Deep Submodular Peripteral Networks

### v2

@@ 3 Peripteral Loss
In this section, we describe a new “peripteral” loss to train a DSPN and simultaneously address how to learn, in a modern machine learning context, from numerically graded pairwise comparisons (GPCs) queryable from a target oracle.
@@
On the other hand, a ratio alone is not a good objective to maximize, one big reason being that small changes in the denominator can have disproportionately large effects on the value of the ratio, which can amplify noise and lead to erratic optimization behavior, not to mention ${\delta}/0$ issues [^2] In the sequel, we use ${\Delta}$ and ${\delta}$ where the $E,M$ pair is implicit. which we address further below. Ideally, we could find a hinge-like loss that will penalize small ratios and reward large ratios. Consider the following:
@@
Here, if ${\Delta}$ is (very) positive, then ${\delta}$ must also be (very) positive to produce a small loss. If ${\Delta}$ is (very) negative then so also must ${\delta}$ be to produce a small loss. The absolute value of ${\Delta}$ determines how non-smooth the loss is so we include an outer pre-multiplication by $|{\Delta}|$ to ensure the slope of the loss, as a function of ${\delta}$, asymptotes to negative one (or one when ${\Delta}$ is negative) in the penalty region, analogous to hinge loss, and thus produces numerically stable gradients. Also analogous to hinge, we add an extra “1” within the $\exp()$ function as a form of margin confidence, analogous to hinge loss that, as with SVMs, expresses a distance between the decision boundary and the support vectors. Plots of $\mathcal{L}_{\Delta}({\delta})$ are shown in Fig. 7 in Appendix F which also discusses the smoothness of $\mathcal{L}_{\Delta}({\delta})$ via its gradients.
Having established the “mother” loss $\mathcal{L}_{\Delta}({\delta})$ we augment it with additional hyperparameters in way that produces smoother and more stable and convergent optimization. Firstly, rather than a requiring a margin of one, we introduce a hyperparameter $\tau$ to express this margin whose default value is $\tau=1$. Second, since the units of the oracle target might be different than the units of the learner, we introduce a unit adjustment parameter $\kappa$ that scales ${\Delta}$. We also introduce an anti-smoothness parameter $\beta>0$ (smaller $\beta$ means smoother loss) to further optionally smooth the transition between the linearly increasing loss region and the zero-loss region of ${\delta}$. Lastly, if ${\Delta}$ gets small, to avoid finite-precision numerical problems, we add a small signed value $\epsilon\operatorname{sgn}({\Delta})$ for $\varepsilon\geq 0$ which slightly pushes the denominator away from zero depending on the sign of ${\Delta}$. This yields $\mathcal{L}_{{\Delta};\beta,\tau,\kappa,\varepsilon}({\delta})$ which is defined as:
The rate of the gating is determined by $\alpha\geq 0$, and since $\frac{\alpha}{|\kappa{\Delta}|}\geq 0$, the gate switches to the absolute value $|{\delta}|$ whenever ${\Delta}$ gets very small. Thus if the oracle target is indifferent to $E$ vs. $M$ this encourages the learner also to be indifferent. Despite these hyperparameters, the loss gradient is still independent of the margin and $\beta$ making it useful for fine-tuning a learning rate (see §F) for gradient-based learning.
Comparing with standard contrastive losses and deep metric learning (Balestriero et al., 2023), a peripteral loss on ${{\delta}_{{f_{w}}}({E}|{M})}=f_{w}(E)-f_{w}(M)$ can be seen as a high-order self-normalizing by $\Delta$ generalization of triplet loss (Weinberger and Saul, 2009) if $E$ is a heterogeneous and $M$ a homogeneous set of objects. We recover aspects of a triplet loss if we set $E=\{v,v^{-}\}$ and $M=\{v,v^{+}\}$ where $M$ is the positive pair (close and homogeneous) while $E$ is the negative pair (distant from each other and thus a diverse pair). With $|E|>2$ and $|M|>2$, we naturally represent high order relationships amongst both the positive group $M$ and the disperse group $E$. We remind the reader that while our peripteral loss could be used for GPC-style contrastive training for representation learning or for RLHF, our experiments in this paper in §5 focus on the submodular learning problem.
@@ 3.1 Augmented Loss for Augmented Data
Since augmented images represent the same information, the submodular gain of augmented images relative to non-augmented images should be small or zero meaning $f_{w}(E|E^{\prime})=f_{w}(E^{\prime}|E)=0$ and $\forall e\in E$, $f_{w}(e|e^{\prime})=f_{w}(e^{\prime}|e)=0$. To encourage this, we use an redundancy regularizer $\mathcal{L}_{\text{redn}}(f_{w};E,E^{\prime})$ defined as follows:
@@ 3.2 Final Loss
With all of the above individual loss components described, the final total loss $\mathcal{L}_{\text{tot}}(f_{w};{{\Delta}({E}|{M})},{{\delta}_{{f_{w}}}({E}|{M})})$ for an $E,M$ pair becomes:
@@ 4 Sampling (E,M) Pairs
Training a DSPN via the peripteral loss requires a dataset $\mathcal{D}=\{(E_{i},M_{i})\}_{i=1}^{N}$ of pairs of sets to be contrasted with each other where each $E_{i},M_{i}\subseteq V$. Since the DSPN starts with the pillar stage, the learnt model can then be tested on samples that are outside of $V$ as we do in §5. Still, exhaustively training on all pairs of subsets from $V$ is intractable. Therefore, the pairing strategy used to incorporate the sampled sets is a critical component in our dataset construction. We thus propose several practical, both passive and active, sampling strategies to efficiently sample from this combinatorial space. The passive strategies are heuristic-based, while the active strategies are dependent on either the DSPN as it is being learnt or optionally the target function if it is possible to optimize. While we describe an outline of these strategies here, appendix E offers further discussion and analysis.
@@
Passive Sampling: If we assume the maximum set size to be $K$ so that $|E|\leq K$ and $|M|\leq K$ and a training dataset to have $C$ classes or clusters, we may generate $E,M$ pairs in two ways. In Style-I, we sample a homogeneous set of size $K$ from a randomly chosen single class and a heterogeneous set of size $K$ randomly over the full dataset. In Style-II, we first choose a random set of $C^{\prime}\leq C$ classes, where $C^{\prime}$ is chosen uniformly at random over $[C]$, and (if $C^{\prime}<C$) subselect from the ground set a set $V_{C^{\prime}}\subseteq V$ containing only those samples corresponding to those $C^{\prime}$ classes. To choose an $E$ set, we perform K-Means clustering on $V_{C^{\prime}}$, and pick the sample closest to each cluster mean to construct our set. To build an $M$ set, we choose one sample for each class/cluster in $V_{C^{\prime}}$ and then add its $\lfloor K/C^{\prime}\rfloor-1$ nearest neighbors. Importantly, the $E$ and $M$ sets in each pair sampled with Style-II are much more difficult to distinguish than Style-I. We also note that these sampling strategies are imperfect, so the supposedly heterogeneous $E$ set may not truly be heterogeneous in relation to its homogeneous counterpart $M$. However, as discussed in §E, the peripteral loss allows us to train the DSPN parameters even if the set pairs are imperfectly labeled since the oracle target just flips the sign of its preference.
@@
Active Sampling: Sets selected by the passive strategies are independent of both the learner and the target function. Thus, due to the combinatorial nature of the problem, there may be a discrepancy between what sets are valued highly by the DSPN and/or the target. This leads us to employ actively sampled sets obtained through two approaches. We refer to the first approach as DSPN Feedback, which obtains sets by maximizing or minimizing the DSPN as it is being learnt. As DSPN is submodular, we may optimize it over either full ground set $V$ or a class-restricted ground set to get high-valued subsets, where the greedy algorithm can be used for maximization (Nemhauser et al., 1978) and submodular minimization (Fujishige, 2005) can be used to find the homogeneous sets (although in practice, submodular minimization heuristics suffice). We refer to the second approach as Target Feedback, where we obtain sets by maximizing the target. The latter approach is only applicable when the target is optimizable (e.g., submodular), which would not be the case with a human oracle. While we do test “Target Feedback” below, our experiments in §5.3 show that it is not necessary for good performance, which portends well for our framework to be used in other GPC contexts (e.g., RLHF reward learning). Thus, while the $E,M$ pair generation strategies mentioned here are specific to DSPN learning, our peripteral loss could be used to train any deep set or other neural network given access to graded oracle preferences, thus yielding a general purpose GPC-style learning procedure.
We now provide a 2D visual example of what passive sets look like in a simplified setting, shown in fig. 2 as well as illustrate the method in which sets are paired in fig. 2. Lastly, in fig. 3 we provide the structure of a DSPN as well as the control flow of how a DSPN is trained; please refer to appendix G for more details.
@@
Figure 2: Set Pairing. We actively sample sets by optimizing the DSPN as it is being learnt or optionally the target function. The depicted graph demonstrates how the actively sampled sets are integrated into $\mathcal{D}=\{(E_{i},M_{i})\}_{i}^{N}$. Th red/blue vertices refer to passively/actively sampled sets. An edge between two vertices indicates that sets from the categories denoted by the vertices are used to create $(E,M)$ pairs to train the DSPN.
Figure 3: The structure of a DSPN as well as the control flow of how a DSPN is trained; parameters are shared between the DSPNs processing E and M sets.
@@ 5 Experiments
Figure 4: Transfer. We evaluate the target function on sets generated by running greedy max on DSPN’s trained with various losses, as well as randomly generated sets. We demonstrate that the DSPN trained with the peripteral loss generates sets that achieve the highest normalized FL evaluations across budgets and datasets.
Figure 5: Offline Experimental Design. We compare different summarization procedures, and assess them based on the test accuracy that a linear model attains upon being trained on the summary. We find that maximizing the learnt DSPN generates high quality datasets, outperforming existing summarization techniques such as CRAIG and k-centers. In many cases, the DSPN approaches the Target FL even though the latter is aware of the duplicates present in the data.
Figure 4: Transfer. We evaluate the target function on sets generated by running greedy max on DSPN’s trained with various losses, as well as randomly generated sets. We demonstrate that the DSPN trained with the peripteral loss generates sets that achieve the highest normalized FL evaluations across budgets and datasets.
Figure 5: Offline Experimental Design. We compare different summarization procedures, and assess them based on the test accuracy that a linear model attains upon being trained on the summary. We find that maximizing the learnt DSPN generates high quality datasets, outperforming existing summarization techniques such as CRAIG and k-centers. In many cases, the DSPN approaches the Target FL even though the latter is aware of the duplicates present in the data.
Figure 6: Online Experimental Design. The performance of a linear probe, trained based on a subset selected by an algorithm that does not use labels, is reported at varying budgets. DSPN-online achieves far higher performance than other streaming algorithms such as reservoir sampling and VeSSAL and is performs comparably to offline algorithms.
@@
We evaluate the effectiveness of our framework by learning a DSPN that can emulate a costly target oracle facility location function. We assert that DSPN training is deemed successful if the subsets recovered by maximizing the learnt DSPN are (1) assigned high values by the target function (§5.1); (2) are usable for some downstream task such as training a predictive model (§5.2). See §H for hyperparameter details.
@@
Datasets: We consider four image classification datasets: Imagenette, Imagewoof, CIFAR100, and Imagenet100. Imagenette, Imagewoof, and Imagenet100 are subsets of Imagenet-1k (Russakovsky et al., 2015), where a subset of the original classes is used. Unlike Imagenet100/CIFAR100, Imagenette and Imagewoof have 10 classes.
@@
DSPN Architecture: We use CLIP ViT encoder (Radford et al., 2021) for all datasets to get the embeddings that are used as inputs to the DSPN pillar. We use a 2-layer network as the pillar (${\phi_{{w_{\mathfrak{p}}}}}$); for DSPN roof (${\psi_{{w_{\mathfrak{r}}}}}$) we use a 2-layered DSF for Imagenette/Imagewoof and 1-layered DSF for CIFAR100/Imagenet100. We use multiple concave activation functions, i.e., all of $\sigma(x)\in\{x,x^{0.5},\ln(1+x),\tanh(x),1-e^{-x}\}$ in every layer of the roof (ensuring $w_{\mathfrak{r}}\geq 0$ during training) and thus allow the training procedure to decide between them.
@@
Training and Evaluation: We train up to sets of size $|E|,|M|\leq 100$ for all datasets. However, we show that we generalize beyond this size even on held-out data. Given ground set $V$, we construct pairs $(E,M)$ both passively and actively, as discussed earlier. Please refer to §G for the details on training. For the target oracle, we use a facility location function with similarity instantiated using a radial basis function kernel with a tuned kernel width. During the evaluation, both learned set functions and target oracle use a held-out ground set $V^{\prime}$ of images, that is, $V^{\prime}\cap V=\emptyset$.
@@ 5.1 Transfer from Target to Learner
We demonstrate that the target FL assigns high values to summaries that are generated by the DSPN in Figure 6. Specifically, we determine if $S$ is a good quality set if it has a high normalized FL evaluation $\frac{f_{\mathfrak{t}}(S)}{f_{\mathfrak{t}}(S_{FL})}$, where $f_{\mathfrak{t}}$ is the target FL and $S_{\mathfrak{t}}$ is the set generated by applying cardinality constrained greedy maximization to the target FL. Therefore, we compare different DSPN models based on the normalized FL evaluation that their maximizers attain. We compare DSPN models trained with three different losses: the peripteral loss, the contrastive max-margin style structured prediction loss (Taskar et al., 2005), and a regression loss (shown as DSPN (peripteral), DSPN (margin) and DSPN (regression) respectively in Figure 6). Please refer appendix H for the definition of baseline losses (Equations 28 and 29). We also compare against randomly generated sets, which are denoted as Random. From this set of experiments, we find that the DSPN trained with the peripteral loss consistently outperforms all other baselines across different summary sizes (budgets) and datasets. Interestingly, the DSPN outperforms random even on budgets greater than 100, despite never having encountered such subsets during training. This provides evidence that the peripteral loss is the loss of choice for training DSPNs.

### v3

@@ 3 Peripteral Loss
We next describe a new “peripteral” loss to train a DSPN and simultaneously address how to learn from numerically graded pairwise comparisons (GPCs) queryable from a target oracle.
@@
On the other hand, a ratio alone is not a good objective to maximize, one big reason being that small changes in the denominator can have disproportionately large effects on the value of the ratio, which can amplify noise and lead to erratic optimization behavior. [^2] In the sequel, we use ${\Delta}$ and ${\delta}$ where the $E,M$ pair is implicit. Ideally, we could find a hinge-like loss that will penalize small ratios and reward large ratios. Consider the following:
@@
Here, if ${\Delta}$ is (very) positive, then ${\delta}$ must also be (very) positive to produce a small loss. If ${\Delta}$ is (very) negative then ${\delta}$ also must be small to produce a small loss. The absolute value of ${\Delta}$ determines how non-smooth the ratio is so we include an outer pre-multiplication by $|{\Delta}|$ to ensure the slope of the loss, as a function of ${\delta}$, asymptotes to negative one (or one when ${\Delta}$ is negative) in the penalty region, analogous to hinge loss, and thus produces numerically stable gradients. Also analogous to hinge, we add an extra “1” within the $\exp()$ function as a form of margin confidence, analogous to hinge loss that, as with SVMs, expresses a distance between the decision boundary and the support vectors. Plots of $\mathcal{L}_{\Delta}({\delta})$ are shown in Fig. 7 in Appendix F — this appendix also discusses the smoothness of $\mathcal{L}_{\Delta}({\delta})$ via its gradients, describes a few additional useful hyperparameters, and provides a numerically stable solution to the case when $\Delta=0$ (thereby solving any ${\delta}/0$ issues).
Comparing with standard contrastive losses and deep metric learning [8], a peripteral loss on ${{\delta}_{{f_{w}}}({E}|{M})}=f_{w}(E)-f_{w}(M)$ can be seen as a high-order self-normalizing by $\Delta$ generalization of triplet loss [108] if $E$ is a heterogeneous and $M$ a homogeneous set of objects. We recover aspects of a triplet loss if we set $E=\{v,v^{-}\}$ and $M=\{v,v^{+}\}$ where $M$ is the positive pair (close and homogeneous) while $E$ is the negative pair (distant from each other and thus a diverse pair). With $|E|>2$ and $|M|>2$, we naturally represent high order relationships amongst both the positive group $M$ and the disperse group $E$ (which is apparently rare, see§A.1). We remind the reader that while our peripteral loss could be used for GPC-style contrastive training for representation learning or for RLHF-based transformer alignment, our experiments in this paper in §5 focus on the submodular learning problem.
@@ 3.1 Augmented Loss for Augmented Data
Since augmented images represent the same information, the submodular gain of augmented images relative to non-augmented images should be small or zero meaning $f_{w}(E|E^{\prime})=f_{w}(E^{\prime}|E)=0$ and $\forall e\in E$, $f_{w}(e|e^{\prime})=f_{w}(e^{\prime}|e)=0$. To encourage this, we use a redundancy regularizer $\mathcal{L}_{\text{redn}}(f_{w};E,E^{\prime})$ defined as follows:
@@ 3.2 Final Loss
With all of the above individual loss components now well described, our final total loss $\mathcal{L}_{\text{tot}}(f_{w};{{\Delta}({E}|{M})},{{\delta}_{{f_{w}}}({E}|{M})})$ for an $E,M$ pair becomes:
$\mathcal{L}_{\text{tot}}(f_{w};{{\Delta}({E}|{M})},{{\delta}_{{f_{w}}}({E}|{M})})$ $=\mathcal{L}_{{{\Delta}({E}|{M})}}({{\delta}_{{f_{w}}}({E}|{M})})+\mathcal{L}_{\text{aug}}(f_{w};E\cup M,E^{\prime}\cup M^{\prime})$
@@
Fig. 1 shows the structure of a DSPN and its learning control flow. Please see appendix G for more details.
Figure 1: The structure of a DSPN and the control flow of how a DSPN is trained; parameters are shared between the DSPNs processing E and M sets.
@@ 4 Sampling (E,M) Pairs
DSPN training via the peripteral loss requires a dataset $\mathcal{D}=\{(E_{i},M_{i})\}_{i=1}^{N}$ of pairs of sets where each $E_{i},M_{i}\subseteq V$. Since the DSPN starts with the pillar stage, the learnt model can then be tested on samples that are outside of $V$ as we do in §5. Still, exhaustively training on all pairs of subsets from $V$ is infeasible. Therefore, the pairing strategy used to incorporate the sampled sets is a critical component in our dataset construction. We thus propose several practical, both passive and active, strategies to efficiently sample from this combinatorial space. The passive strategies are heuristic-based, while the active strategies are dependent on either the DSPN as it is being learnt or optionally the target function if it is possible to optimize (fortunately, Table 1 shows target sampling is not necessary). Our description here of these strategies is augmented with further discussion and analysis in appendix E.
@@
Passive Sampling: If we assume the maximum set size to be $K$ so that $|E|\leq K$ and $|M|\leq K$ and a training dataset to have $C$ classes or clusters, we may generate $E,M$ pairs in two ways. In Style-I, we sample a homogeneous set of size $K$ from a randomly chosen single class and a heterogeneous set of size $K$ randomly over the full dataset. In Style-II, we first choose a random set of $C^{\prime}\leq C$ classes, where $C^{\prime}$ is chosen uniformly at random over $[C]$, and (if $C^{\prime}<C$) subselect from the ground set a set $V_{C^{\prime}}\subseteq V$ containing only those samples corresponding to those $C^{\prime}$ classes. To choose an $E$ set, we perform K-Means clustering on $V_{C^{\prime}}$, and pick the sample closest to each cluster mean to construct our set. To build an $M$ set, we choose one sample for each class/cluster in $V_{C^{\prime}}$ and then add its $\lfloor K/C^{\prime}\rfloor-1$ nearest neighbors. Importantly, the $E$ and $M$ sets in each pair sampled with Style-II are much more difficult to distinguish than Style-I. We also note that these sampling strategies are imperfect, so the supposedly heterogeneous $E$ set may not truly be heterogeneous in relation to its homogeneous counterpart $M$. However, as discussed in §E, the peripteral loss allows us to train a DSPN even if the set pairs are imperfectly labeled since the oracle just flips the sign of its preference. Simplified passive 2D sets are shown in Fig. 3 and a set-pairing illustration is seen in Fig. 3.
@@
Active Sampling: Sets selected by the passive strategies are independent of both the learner and the target function. Thus, due to the combinatorial nature of the problem, there may be a discrepancy between what sets are valued highly by the DSPN and/or the target. This leads us to employ actively sampled sets obtained through two approaches. We refer to the first approach as DSPN Feedback, which obtains sets by maximizing or minimizing the DSPN as it is being learnt. As DSPN is submodular, we may optimize it over either full ground set $V$ or a class-restricted ground set to get high-valued subsets, where the greedy algorithm can be used for maximization [73] and submodular minimization [28] can be used to find the homogeneous sets (although in practice, submodular minimization heuristics suffice). We refer to the second approach as Target Feedback, where we obtain sets by optimizing the target. The latter approach is only applicable when the target is optimizable (e.g., submodular), which would not be the case with a human oracle. While we do test “Target Feedback” below, our experiments in §5.3 show that it is not necessary for good performance, which portends well for our framework to be used in other GPC contexts (e.g., RLHF learning via human feedback). Thus, while the $E,M$ pair generation strategies mentioned here are specific to DSPN learning, our peripteral loss could be used to train any deep set or other neural network given access to graded oracle preferences, thus yielding a general purpose GPC-style learning procedure.
@@
Figure 3: Set Pairing. We actively sample sets by optimizing the DSPN as it is being learnt or optionally the target function. The depicted graph demonstrates how the actively sampled sets are integrated into $\mathcal{D}=\{(E_{i},M_{i})\}_{i}^{N}$. Th red/blue vertices refer to passively/actively sampled sets. An edge between two vertices indicates that sets from the categories denoted by the vertices are used to create $(E,M)$ pairs to train the DSPN. Dashed edges represent non-critical pairs.
@@ 5 Experiments
We evaluate the effectiveness of our framework by training a DSPN to emulate a costly target oracle FL function. We assert that DSPN training is deemed successful if the subsets recovered by maximizing the learnt DSPN are (1) assigned high values by the target function (§5.1); and (2) are usable for a real downstream task such as training a predictive model (§5.2).
@@
Datasets: We consider four image classification datasets: Imagenette, Imagewoof, CIFAR100, and Imagenet100. Imagenette, Imagewoof, and Imagenet100 are subsets of Imagenet-1k [84] where a subset of the original classes is used. Imagenet100 (resp. CIFAR100) have 100 classes and 130,000 (resp. 60,000) images, while Imagenette and Imagewoof have 10 classes (and 13,000 images).
@@
DSPN Architecture: We use CLIP ViT encoder [82] for all datasets to get the embeddings that are used as inputs to the DSPN which can be seen as a fixed part of the pillar. We use an additional 2-layer network as the trainable part of the pillar (${\phi_{{w_{\mathfrak{p}}}}}$); for DSPN roof (${\psi_{{w_{\mathfrak{r}}}}}$) we use a 2-layered DSF for Imagenette/Imagewoof and 1-layered DSF for CIFAR100/Imagenet100. We use multiple concave activation functions, i.e., all of $\sigma(x)\in\{x,x^{0.5},\ln(1+x),\tanh(x),1-e^{-x}\}$ in every layer of the roof (ensuring $w_{\mathfrak{r}}\geq 0$ during training) and thus allow the training procedure to decide between them. See §H for hyperparameter details.
Target: We use Facility Location (FL) as a target whose construction is rather expensive (quadratic) and also does not generalize to held-out data, but this is a key point. That is, we show how we distill from this expensive target FL to a cheaper and generalizable DSPN. Each target FL is created by computing a real-valued feature vector for each sample and then computing a non-negative similarity between each such vector. We used a CLIP ViT encoder to encode input images, and a tuned RBF kernel to construct similarities for our target FLs.
@@
Training and Evaluation: We train up to sets of size $|E|,|M|\leq 100$ for all datasets. We show that we generalize beyond this size not just on held-out subsets but even on held-out data. Given ground set $V$, we construct pairs $(E,M)$ both passively and actively, as discussed earlier. Please refer to §G for the details on training. For the target oracle, we use a facility location function with similarity instantiated using a radial basis function kernel with a tuned kernel width. During the evaluation, both learned set functions and target oracle use a held-out ground set $V^{\prime}$ of images, that is, $V^{\prime}\cap V=\emptyset$.
@@ 5.1 Transfer from Target to Learner
We demonstrate that the target FL assigns high values to summaries that are generated by the DSPN in Figure 4. Specifically, we determine if $S$ is a good quality set if it has a high normalized FL evaluation $\frac{f_{\mathfrak{t}}(S)}{f_{\mathfrak{t}}(S_{FL})}$, where $f_{\mathfrak{t}}$ is the target FL and $S_{\mathfrak{t}}$ is the set generated by applying cardinality constrained greedy maximization to the target FL.
Figure 4: Transfer. We compare different loss functions in terms of their effectiveness at training a DSPN.
Therefore, we compare different DSPN models based on the normalized FL evaluation that their maximizers attain. We compare DSPN models trained with three different losses: the peripteral loss, the contrastive max-margin style structured prediction loss [98], and a regression loss (shown as DSPN (peripteral), DSPN (margin) and DSPN (regression) respectively in Figure 4). Please refer appendix H for the definition of baseline losses (Equations 30 and 31). We also compare against randomly generated sets, which are denoted as Random. From this set of experiments, we find that the DSPN trained with the peripteral loss consistently outperforms all other baselines across different summary sizes (budgets) and datasets. Interestingly, the DSPN outperforms random even on budgets greater than 100, despite never having encountered such subsets during training. This provides evidence that our GPC-based peripteral loss is the loss of choice for training DSPNs.
Figure 5: Offline Experimental Design. We compare different summarization procedures, and assess them based on the test accuracy that a linear model attains upon being trained on the summary. We find that maximizing the learnt DSPN generates high quality datasets, outperforming existing summarization techniques such as CRAIG and k-centers. In many cases, the DSPN approaches the Target FL even though the latter is aware of the duplicates present in the data. Takeaway: DSPN effectively chooses training samples for labeling from an unlabeled pool.
Figure 6: Online Experimental Design. The performance of a linear probe, trained based on a subset selected by an algorithm that does not use labels, is reported at varying budgets. DSPN-online achieves far higher performance than other streaming algorithms such as reservoir sampling and VeSSAL and is performs comparably to offline algorithms. Takeaway: DSPN can effectively select training samples for labeling from a stream of unlabeled data.